# Feature Attribution with Necessity and Sufficiency via Dual-stage Perturbation Test for Causal Explanation

### v1

@@ 4 Feature Attribution as a Problem of PNS Measurement
Therefore, in order to calculate the target score as shown in Eq. 6, it can be broken down into the calculation of four probabilities, where $P(A_{\mathbf{s}},B_{\mathbf{s}})$ and $P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})$ are the observation probabilities, and $P(\bar{B}_{\bar{A}_{\mathbf{s}}}|A_{\mathbf{s}},B_{\mathbf{s}})$ and $P(B_{{\mathbf{s}}_{A_{\mathbf{s}}}}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})$ are the counterfactual probabilities PN and PS, respectively. To facilitate calculation, we further formalize PN and PS into the following forms.
$P(\bar{B}_{{\mathbf{s}}_{\bar{A}_{\mathbf{s}}}}|A_{\mathbf{s}},B_{\mathbf{s}})$ (7)
$=\mathbb{E}_{\mathbf{x}\sim P(\tilde{\mathbf{X}}|A_{\mathbf{s}},B_{\mathbf{s}})}[P(|f(g(\mathbf{x},\bar{\mathbf{s}},\mathbf{x}^{\prime}))-f(\mathbf{x})|\leq\!c)_{~{}\!\text{do}(\tilde{\mathbf{X}}=\mathbf{x})})],$
$P(B_{{\mathbf{s}}_{A_{\mathbf{s}}}}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})$ (8)
$=\mathbb{E}_{\mathbf{x}\sim P(\tilde{\mathbf{X}}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})}[P(|f(g(\mathbf{x},{\mathbf{s}},\mathbf{x}^{\prime}))-f(\mathbf{x})|>\!c)_{~{}\!\text{do}(\tilde{\mathbf{X}}=\mathbf{x})})],$
where $\text{do}(\tilde{\mathbf{X}}$$=$$\mathbf{x})$ represents an intervention on variable $\tilde{\mathbf{X}}$ that fixes the value as $\mathbf{x}$. Note that event $A$ not occurring is approximated by the complementary feature set $\bar{\mathbf{s}}$. $P(\bar{B}_{\bar{A}_{\mathbf{s}}}|A_{\mathbf{s}},B_{\mathbf{s}})$ and $P(B_{{\mathbf{s}}_{A_{\mathbf{s}}}}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})$ will be calculated by the necessity and sufficiency modules in Section 5.
5 Necessary and Sufficient Attribution Estimation via Dual-stage Perturbation Test
@@ 5.1 Overview
Based on the definitions of necessary and sufficient attribution, we propose a Feature Attribution with Necessity and Sufficiency (FANS) method to extract the feature subset with the highest PNS, as shown in Figure 2. FANS consists of necessity modules and sufficiency modules, each performing our proposed dual-stage (factual and interventional stage) perturbation test to estimate PN and PS respectively. Since the conditional distributions involved in the factual stage may be complex, FANS adopts the sampling-importance-resampling (SIR) method to approximate these distributions with observed samples. Finally, FANS integrates the outputs of both modules into PNS and optimizes the perturbation. FANS assumes that the function $f$ is a classification model with the probability of each class as output.
5.2 Dual-stage Perturbation Test
We design two different dual-stage perturbation tests to estimate the PS in Eq. 8 and PN in Eq. 7, respectively, by following the Abduction-Action-Prediction counterfactual reasoning paradigm (Pearl, 2013). Abduction involves setting conditions for specific variables (corresponds to $\tilde{\mathbf{X}}$ in our context) based on the observations (corresponds to prediction $\mathbf{Y}$ given a perturbation $\mathbf{S}$). The action involves intervening in the values of certain variables (corresponds to $\mathbf{S},\tilde{\mathbf{X}}$). Prediction involves computing the resulting conditional predictive distribution. This corresponds to the conditional distribution for $\mathbf{Y}$ given intervention value of $\mathbf{S}$ and the values of $\tilde{\mathbf{X}}$ that satisfy the factual observations.
We first present the dual-stage perturbation test for PS. 1) (Factual stage) Draw inputs from a distribution conditional on the fact that predictions remain unchanged after using the perturbation function (Eq. 2) to perturb the features of $\tilde{\mathbf{X}}$ on $\overline{\mathbf{s}}$. 2) (Interventional stage) Remove all arrows pointing to $\tilde{\mathbf{\mathbf{X}}}$ in Figure 1 (i.e., perform the “do” operator in causality). For each collected sample on the feature $\mathbf{s}$, use the perturbation function in Eq. 2 to randomly perturb multiple times and calculate the proportion of prediction changes. Finally, average the proportion of prediction changes for each input.
Similarly, the dual-stage perturbation test for PN is constructed by exchanging the perturbations and prediction events involved in the above two stages respectively.
5.3 Sufficiency Module
In this part, we provide a description of how to estimate probability of sufficiency with the help of our sufficiency module.
5.3.1 Factual Stage
[…]
$\!w_{\text{SF}}(\mathbf{x})\!=\!\left\{\begin{array}[]{lr}\!\!\!r\!\cdot\!P(|f(g(\mathbf{x},\bar{\mathbf{s}},\mathbf{x}^{\prime}))\!-\!f(\mathbf{x})|\!\leq c|~{}\mathbf{x}),&\!\!\!\!\!\!\!\text{ if }\|\mathbf{x}_{\bar{\mathbf{s}}}\!-\!\mathbf{x^{\prime}}_{\bar{\mathbf{s}}}\|_{p}\leq b\\ \!\!\!0,&\text{ otherwise.}\end{array}\right.$ (9)
Hence, the resampling weight of $\mathbf{x}$ is determined by two conditions: $\|\mathbf{x}_{\mathbf{s}}-\mathbf{x^{\prime}}_{\mathbf{s}}\|_{p}\leq b$ and $|f(g(\mathbf{x},\bar{\mathbf{s}},\mathbf{x^{\prime}}))\!-\!f(\mathbf{x})|\!\leq c$. To enhance the smoothness of the weight output from $w_{\text{SF}}(\cdot)$, we introduce the Gaussian kernel function to soften the above two conditions and take the product of their outputs as an approximation of the resampling weight $w_{\text{SF}}(\mathbf{x})$ of $\mathbf{x}\sim P(\mathbf{X})$, as follows.
$\small\tilde{w}_{\text{SF}}(\mathbf{x}):=\exp\big{(}\frac{\|\mathbf{x}_{\bar{\mathbf{s}}}-\mathbf{x^{\prime}}_{\bar{\mathbf{s}}}\|^{2}_{p}}{-2b^{2}}\big{)}\cdot\exp\big{(}\frac{|f(g(\mathbf{x},\bar{\mathbf{s}},\mathbf{x^{\prime}}))\!-\!f(\mathbf{x})|^{2}}{-2c^{2}}\big{)},$ (10)
where the thresholds $b,c$ are used as the bandwidths of the Gaussian kernel. Further, we feed observed samples and their weight according to Eq. 12 into the SIR algorithm, and the sample set that approximates the conditional distribution $P(\tilde{\mathbf{X}}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})$ in PS (Eq. 8) can be obtained.
5.3.2 Intervention Stage
Let $\mathcal{E}_{\text{SF}}$ be a sample set obtained through SIR in the factual stage of our Sufficiency Module. Then, PS (Eq. 8) can be estimated as follows:
$\hat{\text{PS}}=\frac{1}{|\mathcal{E}_{\text{SF}}|\cdot t}\sum_{\mathbf{x}\in\mathbf{S}_{\text{SF}}}\sum^{t}_{j=1}\mathbb{I}(|f(g({\mathbf{x}},\mathbf{s},\mathbf{x}^{\prime}))-f({\mathbf{x}})|>c),$ (11)
where the constant $t$ denotes the number of perturbations.
5.4 Necessity Module
[…]
$\small\!\!\tilde{w}_{\text{NC}}(\mathbf{x})\!:=\!\exp\big{(}\frac{\|\mathbf{x}_{\mathbf{s}}\!-\!\mathbf{x^{\prime}}_{\mathbf{s}}\|^{2}_{p}}{-2b^{2}}\big{)}\cdot\big{(}1\!-\exp\big{(}\frac{|f(g(\mathbf{x},\mathbf{s},\mathbf{x}^{\prime}))\!-\!f(\mathbf{x})|^{2}}{-2c^{2}}\big{)}\big{)}.$ (12)
Further, let $\mathcal{E}_{\text{NC}}$ be a sample set obtained through SIR in the factual stage of Necessity Module. Then PN (Eq. 7) can be estimated as follows:
$\small\hat{\text{PN}}=\frac{1}{|\mathcal{E}_{\text{NC}}|\cdot u}\sum_{\mathbf{x}\in\mathcal{E}_{\text{NC}}}\sum^{u}_{j=1}\mathbb{I}(|f(g(\mathbf{x},\mathbf{s},\mathbf{x}^{\prime}))-f(\mathbf{x})|\leq c),$ (13)
where the constant $u$ denotes the number of perturbations.
5.5 Optimization
In this section, we aim to maximize the estimated PNS with respect to perturbation $\mathbf{s}$ by utilizing gradient-based optimization algorithms.
As mentioned in Section 4, since probability $P(A_{\mathbf{s}},B_{\mathbf{s}})$ (or $P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})$) can be expressed as the weighted sum of conditional probabilities $P(A_{\mathbf{s}},B_{\mathbf{s}}|\mathbf{x})$ with respect to all possible $\mathbf{x}$, where the weight is $P(\mathbf{X})$, we can estimate $P(A_{\mathbf{s}},B_{\mathbf{s}})$ and $P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})$ by summing their SIR-based weights of the observed samples in Eq. 12 and Eq. 10 respectively, as follows. Given the sample set $\mathcal{E}$ that are independently drawn from the distribution $P$($\mathbf{X}$),
$\small\!\!\hat{P}(A_{\mathbf{s}},B_{\mathbf{s}})=\sum_{\mathbf{x}\in\mathcal{E}}w_{\text{NC}}(\mathbf{x}),\quad\hat{P}(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})=\sum_{\mathbf{x}\in\mathcal{E}}w_{\text{SF}}(\mathbf{x})$ (14)
@@ 5.5 Optimization
Second, to optimize Eq. 5 with respect to perturbation $\mathbf{s}$, we can use a gradient-based algorithm (e.g., stochastic gradient descent) for optimization. Given the discrete nature of the set s, we can first encode it as a binary vector containing d dimensions, and then relax the entries in the discrete set $\{0,1\}$ into the continuous range $[0,1]$. Additionally, we use $\bm{1}$--$\mathbf{s}$ to represent the complement of $\mathbf{s}$, where $\bm{1}$ is an all-ones vector. Further, we replace $\|\mathbf{x}_{\bar{\mathbf{s}}}-\mathbf{x^{\prime}}_{\bar{\mathbf{s}}}\|_{p}$ in Eq. 9 with $\|\mathbf{x}\circ(\bm{1}-\mathbf{s})-\mathbf{x^{\prime}}\circ(\bm{1}-\mathbf{s})\|_{p}$, where $\mathbf{s}$ is now a relaxed binary vector and $\circ$ is a element-wise multiplication. Similarly, we replace $(\bm{1}\!-\!\mathbf{m}_{\mathbf{S}})\circ\tilde{\mathbf{X}}_{\mathbf{S}}+\mathbf{m}_{\mathbf{S}}\circ\mathbf{x}^{\prime}_{\mathbf{S}}$ in Eq. 2 with $(\bm{1}\!-\!\mathbf{m})\circ\tilde{\mathbf{X}}\circ\mathbf{s}+\mathbf{m}\circ\mathbf{x}^{\prime}\circ\mathbf{s}$. In addition, the indicator function $\mathbb{I}(|f(\cdot)\!-\!f(\cdot)|>c)$ in Eq. 11 is not differentiable, thus we replace it with the smooth function $1\!-\exp{(-|f(\cdot)\!-\!f(\cdot)|)}$ to indicate whether the prediction changes. Similarly, $\mathbb{I}(|f(\cdot)-f(\cdot)|\leq c)$ in Eq. 13 is replaced with $\exp(-|f(\cdot)-f(\cdot)|)$.
Finally, after the optimization with respect to $\mathbf{s}$ is completed by maximizing Eq. 5, the resultant values of $\mathbf{s}$ will be utilized as weights assigned to various features.

### v3

@@ 3 Feature Attribution as a Problem of PNS Measurement
Therefore, to calculate the target score as shown in Eq. 7, it can be broken down into the calculation of the boundary $b$, threshold $c$ and four probabilities: $P(A_{\mathbf{s},b},B_{\mathbf{s},c})$ and $P(\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$ are the observation probabilities, while $P(\bar{B}_{{\mathbf{s},c}_{\bar{A}_{\mathbf{s},b}}}|A_{\mathbf{s},b},B_{\mathbf{s},c})$ and $P(B_{{\mathbf{s},c}_{A_{\mathbf{s},b}}}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$ are the counterfactual probabilities PN and PS, respectively. According to the theory of counterfactual reasoning, PN and PS can be converted into the following equations for calculation (see Appendix A for detailed derivation).
$P(\bar{B}_{{\mathbf{s},c}_{\bar{A}_{\mathbf{s},b}}}|A_{\mathbf{s},b},B_{\mathbf{s},c})$ (8)
$=\mathbb{E}_{\mathbf{x}\sim P(\tilde{\mathbf{X}}|A_{\mathbf{s},b},B_{\mathbf{s},c})}[P(|f(g(\mathbf{x},\bar{\mathbf{s}},\mathbf{x}^{\prime}))-f(\mathbf{x})|\leq\!c)_{~{}\!\text{do}(\tilde{\mathbf{X}}=\mathbf{x})}],$
$P(B_{{\mathbf{s},c}_{A_{\mathbf{s},b}}}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$ (9)
$=\mathbb{E}_{\mathbf{x}\sim P(\tilde{\mathbf{X}}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})}[P(|f(g(\mathbf{x},{\mathbf{s}},\mathbf{x}^{\prime}))-f(\mathbf{x})|>\!c)_{~{}\!\text{do}(\tilde{\mathbf{X}}=\mathbf{x})}],$
where $\text{do}(\tilde{\mathbf{X}}$$=$$\mathbf{x})$ denotes an intervention on $\tilde{\mathbf{X}}$ that fixes the value as $\mathbf{x}$, i.e., perform the do-calculus (Pearl, 2000). To compute the boundary $b$, threshold $c$, and the probabilities $P(\bar{B}_{{\mathbf{s},c}_{\bar{A}_{\mathbf{s},b}}}|A_{\mathbf{s},b},B_{\mathbf{s},c})$ and $P(B_{{\mathbf{s},c}_{A_{\mathbf{s},b}}}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$, we will propose a heuristic strategy for $b,c$ and a novel dual-stage perturbation test which implements the counterfactual reasoning paradigm, as introduced in Section 4.
4 Necessary and Sufficient Attribution
via Dual-stage Perturbation Test
@@ 4.1 Overview
By bridging the gap between PNS and feature attribution, we propose a method called Feature Attribution with Necessity and Sufficiency (FANS) to evaluate the Necessary and Sufficient Attribution (Definition 3.4) of each feature subset of the target input, as illustrated in Figure 2. In particular, we first develop a novel dual-stage (factual and interventional) perturbation test that allows us to estimate counterfactual probabilities of events $A_{\mathbf{s},b},B_{\mathbf{s},c}$. Next, we implement two dual-stage perturbation tests: one for PN and the other for PS, serving as the necessity and sufficiency modules of FANS, respectively. Sequentially, we combine the outputs of these modules to derive PNS. Finally, we propose a heuristic strategy to estimate the boundary $b$ and threshold $c$ that can maximize the PNS of $A_{\mathbf{s},b},B_{\mathbf{s},c}$. Since the conditional distributions involved in the factual stage can be complex, FANS employs the Sampling-Importance-Resampling (SIR) method to approximate these distributions using observed samples. Additionally, we combine FANS with a gradient-based optimization to extract the feature subset with the highest Necessary and Sufficient Attribution.
4.2 Dual-stage Perturbation Test
We design two different dual-stage perturbation tests to estimate the PN (Eq. 8) and PS (Eq. 9) by following the Abduction-Action-Prediction counterfactual reasoning paradigm (Pearl, 2013). Abduction involves setting conditions for specific variables (corresponds to $\tilde{\mathbf{X}}$ in our context) based on observations (corresponds to prediction $\mathbf{Y}$ given the dimensions $\mathbf{S}$). Action involves intervening in the values of certain variables (corresponds to $\mathbf{S},\tilde{\mathbf{X}}$). Prediction involves computing the resulting conditional predictive distribution. This corresponds to the conditional distribution for $\mathbf{Y}$ given the intervention value of $\mathbf{S}$ and the factual values of $\tilde{\mathbf{X}}$.
We first present the dual-stage perturbation test for PS estimation. 1) (Factual stage) Draw inputs from a distribution conditional on the fact that predictions remain unchanged after using the perturbation function (Eq. 3) to perturb the features of $\tilde{\mathbf{X}}$ on $\overline{\mathbf{s}}$ (i.e., events $\bar{A}_{\mathbf{s},b}$ and $\bar{B}_{\mathbf{s},c}$ both occur). 2) (Interventional stage) Remove all arrows pointing to $\tilde{\mathbf{\mathbf{X}}}$ in Figure 1. For each collected sample, use the perturbation function in Eq. 3 to randomly perturb on $\mathbf{s}$ (i.e., event $A_{\mathbf{s},b}$ occur) and calculate the proportion of prediction changes (i.e., event $B_{\mathbf{s},c}$ occur). Finally, average the proportion of prediction changes for each input. Similarly, the dual-stage perturbation test for PN estimation is constructed by exchanging the perturbations and prediction events involved in the above two stages respectively.
4.3 Sufficiency Module
This module provides a concrete implementation of the dual-stage perturbation test for PS estimation.
4.3.1 Factual Stage
To draw samples from $P(\tilde{\mathbf{X}}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$ in Eq. 9, we propose applying SIR on the observed sample set $\mathcal{E}$ that follows $P(\mathbf{X})$ to approximate the required conditional distribution. First, given the subset of dimensions $\bar{\mathbf{s}}$, let $r$ denote a normalization constant and the resampling weight for $\mathbf{x}$ sampled from $P(\mathbf{X})$ is as follows (see Appendix B for derivation).
$\!w_{\text{SF}}(\mathbf{x})\!=\!\left\{\begin{array}[]{lr}\!\!\!r\!\cdot\!P(|f(g(\mathbf{x},\bar{\mathbf{s}},\mathbf{x}^{\prime}))\!-\!f(\mathbf{x})|\!\leq c|~{}\mathbf{x}),&\!\!\!\!\!\!\!\text{ if }\|\mathbf{x}_{\bar{\mathbf{s}}}\!-\!\mathbf{x}^{t}_{\bar{\mathbf{s}}}\|_{p}\leq b\\ \!\!\!0,&\text{ otherwise.}\end{array}\right.$ (10)
Hence, the resampling weight is determined by two conditions: $\|\mathbf{x}_{\bar{\mathbf{s}}}-\mathbf{x}^{t}_{\bar{\mathbf{s}}}\|_{p}\leq b$ and $|f(g(\mathbf{x},\bar{\mathbf{s}},\mathbf{x}^{\prime}))-f(\mathbf{x})|\leq c$. To enhance the smoothness of $w_{\text{SF}}(\cdot)$, we introduce a Gaussian kernel function to soften the above two conditions and take their product as an approximation of $w_{\text{SF}}(\mathbf{x})$, as follows.
$\small\tilde{w}_{\text{SF}}(\mathbf{x}):=\exp\big{(}\frac{\|\mathbf{x}_{\bar{\mathbf{s}}}-\mathbf{x}^{t}_{\bar{\mathbf{s}}}\|^{2}_{p}}{-2b^{2}}\big{)}\cdot\exp\big{(}\frac{|f(g(\mathbf{x},\bar{\mathbf{s}},\mathbf{x}^{\prime}))\!-\!f(\mathbf{x})|^{2}}{-2c^{2}}\big{)},$ (11)
where the methods for estimating $b$ and $c$ will be introduced in Section 4.6. We then feed observed samples and their weight according to Eq. 11 into the SIR algorithm to obtain a sample set that approximates the conditional distribution $P(\tilde{\mathbf{X}}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$ in PS (Eq. 9).
4.3.2 Intervention Stage
Let $\mathcal{E}_{\text{SF}}$ be a sample set of $P(\tilde{\mathbf{X}}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$ obtained through SIR. Then, PS (Eq. 9) can be estimated as follows:
$\hat{\text{PS}}=\frac{1}{|\mathcal{E}_{\text{SF}}|\cdot t_{\text{SF}}}\sum_{\mathbf{x}\in\mathcal{E}_{\text{SF}}}\sum^{t_{\text{SF}}}_{j=1}\mathbb{I}(|f(g({\mathbf{x}},\mathbf{s},\mathbf{x}^{\prime}))-f({\mathbf{x}})|>c),$ (12)
where $t_{\text{SF}}$ is a hyperparameter representing the number of times a random perturbation is applied to $\mathbf{x}$ through the perturbation function $g$ (Eq. 3).
4.4 Necessity Module
The implementation of the necessity module is similar to the sufficiency module, differing only in the opposite perturbation and prediction events $A_{\mathbf{s},b},B_{\mathbf{s},c}$ in each stage. Thus, in the factual stage, the resampling weight for the observed input $\mathbf{x}$ sampled from $P(\mathbf{X})$, which is used to generate the sample set approximating the conditional distribution $P(\tilde{\mathbf{X}}|A_{\mathbf{s},b},B_{\mathbf{s},c})$ in Eq. 8, is given by
$\small\!\!\tilde{w}_{\text{NC}}(\mathbf{x})\!:=\!\exp\big{(}\frac{\|\mathbf{x}_{\mathbf{s}}\!-\!\mathbf{x}^{t}_{\mathbf{s}}\|^{2}_{p}}{-2b^{2}}\big{)}\cdot\big{(}1\!-\exp\big{(}\frac{|f(g(\mathbf{x},\mathbf{s},\mathbf{x}^{\prime}))\!-\!f(\mathbf{x})|^{2}}{-2c^{2}}\big{)}\big{)}.$ (13)
Further, let $\mathcal{E}_{\text{NC}}$ be a sample set of $P(\tilde{\mathbf{X}}|A_{\mathbf{s},b},B_{\mathbf{s},c})$ obtained through SIR. Then PN (Eq. 8) can be estimated as follows:
$\small\hat{\text{PN}}=\frac{1}{|\mathcal{E}_{\text{NC}}|\cdot t_{\text{NC}}}\sum_{\mathbf{x}\in\mathcal{E}_{\text{NC}}}\sum^{t_{\text{NC}}}_{j=1}\mathbb{I}(|f(g(\mathbf{x},\bar{\mathbf{s}},\mathbf{x}^{\prime}))-f(\mathbf{x})|\leq c).$ (14)
where the constant $t_{\text{NC}}$ denotes the number of perturbations.
4.5 PNS Estimation
As indicated by Eq. 6, PNS is equivalent to the weighted sum of PS and PN, with the weights being $P(A_{\mathbf{s},b},B_{\mathbf{s},c})$ and $P(\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$. The weights can be calculated as follows.
@@ 4.5 PNS Estimation
$\small\!\!\hat{P}(A_{\mathbf{s},b},B_{\mathbf{s},c})=\sum_{\mathbf{x}\in\mathcal{E}}w_{\text{NC}}(\mathbf{x}),\quad\hat{P}(\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})=\sum_{\mathbf{x}\in\mathcal{E}}w_{\text{SF}}(\mathbf{x}).$ (15)
4.6 Necessary and Sufficient Attribution Estimation
As illustrated in Definition 3.4, in order to estimate our proposed Necessary and Sufficient Attribution, it is necessary to determine the boundary $b$ and threshold $c$ of events $A_{\mathbf{s},b}$ and $B_{\mathbf{s},c}$ that maximize PNS. FANS employs the following heuristics to compute their values.
The value of $b$ is set to ensure that the neighborhood of $\mathbf{x}^{t}$ is both small and nearly uniformly dense. Thus, the value of $b$ is determined based on the Scott rule (Scott, 1979), a bandwidth estimation method widely used in kernel density estimation, which is given by $b=1.06\cdot|\mathcal{E}|^{\frac{1}{4+d}}$, where $\mathcal{E}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P(\mathbf{X})$ and $d$ is the dimension of the input. On the other hand, the strategy for determining the value of $c$ is as follows. First, we use the normal distribution $N(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ to simulate low-intensity noise, where $\sigma$ is a small constant (set to 0.001 in practice), $\mathbf{0}$ is the all-zero vector, and $\mathbf{I}$ is the identity matrix. Assuming the model’s predictions for inputs with this noise remain unchanged, we define the maximum variance of these predictions as $c:=\max_{\mathbf{x}\in\mathcal{E},\epsilon\in\mathcal{R}}|f(\mathbf{x})-f(\mathbf{x}+\epsilon)|$, where $\mathcal{E}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P(\mathbf{X}),\mathcal{R}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$. The sensitivity analysis of $b$ and $c$ are provided in Appendix E.
4.7 Extracting Feature Subset with the Highest Necessary and Sufficient Attribution
@@
In this section, we aim to efficiently extract the feature subset with the highest estimated Necessary and Sufficient Attribution in the target input. Specifically, we use stochastic gradient descent to maximize the estimated Necessary and Sufficient Attribution with respect to the dimension subset $\mathbf{s}$. Given the discrete nature of the dimension subset s, we first encode it as a binary vector containing $d$ dimensions, and then relax the entries in the discrete set $\{0,1\}$ into the continuous range $[0,1]$. Additionally, we use $\bm{1}$--$\mathbf{s}$ to represent the complement of $\mathbf{s}$, where $\bm{1}$ is an all-one vector. Further, we replace $\|\mathbf{x}_{\bar{\mathbf{s}}}-\mathbf{x}^{t}_{\bar{\mathbf{s}}}\|_{p}$ in Eq. 11 with $\|\mathbf{x}\circ(\bm{1}-\mathbf{s})-\mathbf{x}^{t}\circ(\bm{1}-\mathbf{s})\|_{p}$, where $\mathbf{s}$ is now a relaxed binary vector and $\circ$ represents element-wise multiplication. Similarly, we replace $(\bm{1}\!-\!\mathbf{m}_{\mathbf{S}})\circ\tilde{\mathbf{X}}_{\mathbf{S}}+\mathbf{m}_{\mathbf{S}}\circ\mathbf{x}^{\prime}_{\mathbf{S}}$ in Eq. 3 with $(\bm{1}\!-\!\mathbf{m})\circ\tilde{\mathbf{X}}\circ\mathbf{s}+\mathbf{m}\circ\mathbf{x}^{\prime}\circ\mathbf{s}$. In addition, the indicator function $\mathbb{I}(|f(\cdot)\!-\!f(\cdot)|>c)$ in Eq. 12 is not differentiable, so we replace it with the smooth function $1\!-\exp{(-|f(\cdot)\!-\!f(\cdot)|)}$ to indicate whether the prediction changes. Similarly, $\mathbb{I}(|f(\cdot)-f(\cdot)|\leq c)$ in Eq. 14 is replaced with $\exp(-|f(\cdot)-f(\cdot)|)$.
After optimizing with respect to $\mathbf{s}$ by maximizing Eq. 7, the resultant $\mathbf{s}$ is used for the soft selection of the feature subset with the highest Necessary and Sufficient Attribution.